Suppose the demand for a product is given by D(p)=−3p+139..

A) Calculate the elasticity of demand at a price of $31.

Elasticity = ? (Round to three decimal places.)

B) At what price do you have unit elasticity? (Round your answer to the nearest penny.)

Price = ?
Answer #1

- Given: Demand function D(p)==30+134 @ Price ; p= 31 $ : P = P(31)= -3(31) +134 = 16; : Differentiating © writ. p we get, 46